OpenStudy (anonymous):

When a 58 g tennis ball is served, it accelerates from rest to a constant speed of 36 m/s. The impact with the racket gives the ball a constant acceleration over a distance of 35 cm. What is the magnitude of the net force acting on the ball?

OpenStudy (anonymous):

First find the acceleration using the equation\[v^2=2*a*d\] The force will be F=ma